Suppose that gold and silver are used to settle international balance of payments accounts. If the U.S. dollar is pegged to gold at $36 per ounce and the euro is pegged to gold €54 per ounce and to silver at €4.50 per ounce of silver. If the Japanese yen is pegged to silver at ¥240 per ounce of silver, what is the USD/JPY exchange rate?

Multiple Choice

  • $1 = ¥90

  • None of the options.

  • $1 = ¥110

  • $1 = ¥100

  • $1 = ¥80
Answer #1

Yen/Euro = 240/4.50

i.e. 53.333 Yen/Euro

Also, Euro/Gold = 54/36

i.e. 1.5 Euro/Dollar

Hence, Yen/Dollar = (Yen/Euro)*(Euro/Dollar)

= 53.333*1.5

= 80

Hence, the answer is 1$ = Yen 80

  • $1 = ¥80